# What is the Relationship between Tensor Factorizations and Circuits (and How Can We Exploit it)?

### v1

@@ 7 Empirical Evaluation: Which RG and Layers to use?
Destructuring modern PC architectures (as well as tensor factorizations) into our pipeline (Fig. 8) allows us to create new tensorized architectures by simply following a mix & match approach (Table 1). At the same time, it helps us understand what really matters between different model classes from the point of views of expressiveness, speed of inference and easiness of optimization. We can now in fact easily disentangle key ingredients such as the role of RGs and the choice of composite layers in modern circuit architectures, and pinpoint which is responsible for a boost in performance. For example, HCLTs have been considered as one of the best performing circuit model architectures in recent benchmarks (Liu et al., 2022; 2023a), but until now it has not been clear why they were outperforming other architectures such as RAT-SPNs and EiNets. Within our framework, we can try to answer that question by answering more precise questions: is it the effect of their RG that is learned from data (Section 4.1)?, the use of their composite sum-product layer parameterization (Section 5.1)? or are other hyperparameter choices the culprit? (spoiler: it is going to be the use of CP layers).
Specifically, in this section we are interested in answering the following three research questions following a rigorous empirical investigation. RQ1) What are the computational resources needed (time and GPU memory) at test and training time for some of the many tensorized architectures we can now build? RQ2) What is the impact of the choice of RG and composite sum-product layer on the performance of tensorized circuits trained as distribution estimators? RQ3) Can we retain (most of) the performances of pre-trained tensorized PCs using Tucker layers if we factorize these into CP layers as illustrated in Fig. 12(a) $\rightarrow$ Fig. 12(b)? Note that we are not asking what is the impact of folding (Section 4.4), as we already know the answer: folding is essential for large-scale tensorized architectures. As such, throughout all experiments, we use folded tensorized circuits. We emphasise that the aim of our experiments is not to reach state-of-the-art results in distribution estimation, but rather that of understanding the role of the ingredients of tensorized circuit architectures. All experiments were run on a single NVIDIA RTX A6000 GPU with 48GB of memory.
@@
We will evaluate our architectures by performing distribution estimation on image datasets. We use the Mnist-family, which includes 6 datasets of gray-scale 28x28 images—Mnist (LeCun et al., 2010), FashionMnist (Xiao et al., 2017), and EMNIST with its 4 splits (Cohen et al., 2017)—and the CelebA dataset down-scaled at 64x64 (Liu et al., 2015), which we explore in two versions: one with RGB pixels and the other with pixels preprocessed by the lossless YCoCg color-coding (Malvar & Sullivan, 2003), as recent results suggested that such a transform can greatly lower bpds. [^13] We take this evidence from Liu et al. (2023a; b), which use however a lossy variant of the YCoCg transform that unfortunately artificially inflates likelihoods. As such, their bpds for PCs are not directly comparable with ours, nor with the other deep generative models in their tables. We confirmed this issue in their evaluation protocol via personal communication.
@@
We train circuits to estimate the probability distribution that is assumed to have generated the images, considering each pixel as a random variable. As such, the input units in the circuit represent Categorical distributions having 256 values. For RGB images, we associate three Categorical distribution units per pixel (one per color channel). We perform maximum likelihood by stochastic gradient ascent, i.e., want to maximize the following objective
@@
where $Z=\sum_{\bm{\mathrm{x}}}c(\bm{\mathrm{x}})$ is the partition function of the PC $c$ [^14] After training, one can efficiently “embed” the normalization constant in the parameters of a PC, effectively renormalizing them (and thus yielding a partition function $Z$ equal to $1$), as detailed in Peharz et al. (2015)., and $\mathcal{B}$ a batch of training data. After some preliminary experiments, we found that optimizing PCs with Adam (Kingma & Ba, 2015) using a learning rate of $10^{-2}$ delivered, on average, the best performing models for the datasets we considered. We also settled to reparameterize the circuit sum parameters via clamping and setting $\epsilon=10^{-19}$ (Eq. 11) after each optimization step as to keep them non-negative, as it was giving the best learning dynamics among all possible reparameterizations (Section 3.2). In the following, we will summarize our findings when answering RQ1-3, while distilling recommendations for practitioners on how to build and learn circuits.
@@
For these experiments, we consider the following RGs: PD, as commonly used in architectures such as RAT-SPNs and EiNets, and the two novel light-weight and data-agostic RGs we introduced in Section 4.1, QTs  [^15] Throughout our experiments, we will refer to QT-4 simply as QT. and QGs . We do not consider RND as it is usually just a balanced binary tree (Peharz et al., 2020c), and as such would yield the same time and memory performance of a QT. For the same reason we do not consider CL as they are tree RGs that end up being quasi-balanced after being rooted. [^16] The root is chosen to be the barycenter of the graph to increase parallelism (Dang et al., 2021; 2022c). For layers, we consider Tucker (Eq. Tucker-layer), CP (Eq. CP-layer), CPS (Eq. CPS-layer) and CPXS (Section 5.2).
@@
We emphasise that non-folded versions of these architectures, e.g. RAT-SPNs (Peharz et al., 2020c), can be orders of magnitude slower, therefore hindering both learning and deployment in practice. In Fig. E.1 in the Appendix, we show the results of the same benchmark reported in Fig. 16 but for the CelebA dataset, which is more challenging because it is equivalent to perform distribution estimation on a much higher dimensional space ($12,288=64\times 64\times 3$ instead of $784=28\times 28\times 1$). [^17] Note that for our RQ1, all image datasets with the same resolutions would yield the very same results. So the fact we pick as a grey-scale dataset MNIST and not F-MNIST is not really important.. From this additional experiment, we can conclude that even in higher dimensions the scaling behavior of RGs and layers is the same. Finally, in Fig. E.3, we zoom on a comparison between CPS and CPXS. There, we show that for the same choice of RG and $K$, CPS and CPXS layers require the same time/space resources as expected, with CPXS only being slightly faster at training-time.
Takeaway 1. QT and QG should be preferred to PD as RGs if we want to scale circuits, with the former being the most scalable than the latter. Layer-wise, CP layers scale, as expected, to larger values of $K$ than Tucker layers and for even larger layers parameter sharing (CPS, CPXS) is recommended.
@@
We now test our tensorized architectures as distribution estimators and we consider our mixed&matched architecture from RQ1. For each architecture, we vary the model size by varying $K$, the number of units for each layer, in $\{16,32,64,128,256,512\}$ for the MNIST family and up to $256$ for CelebA. To assess the effect of learning RGs from data, we compare against HCLTs (CL-$\text{CP}^{\top}$in our nomenclature) as reported in (Dang et al., 2022a). For all experiments, we use a batch size of 256, and train for at most 200 epochs stopping training if the validation log-likelihood does not improve after 5 epochs. We use the average test-set bits-per-dimension (bpd) as the evalutation criterion, i.e. $\mathsf{bpd}(\mathcal{D},c)=-\mathcal{L}(\mathcal{D},c)/(d\cdot\log 2)$, where $d$ is the number of features in dataset $\mathcal{D}$ and $\mathcal{L}$ is defined as in Eq. 20.
@@
In Fig. 17 we report the average test-set bpd on Mnist, FashionMnist and CelebA. An immediate visible pattern emerges when comparing the architectures w.r.t. the choice of RG: Both QT- and QG-based architectures outperform those based on PD, and also manage to scale to larger datasets like CelebA. On average, the best performing architectures are those built out of QGs. This is expected as such RGs, different from QTs, allow different partitionings for a same region (and therefore require the usage of mixing layers as discussed in Eq. Mixing-layer). The PD region graphs, despite being DAG-shaped as QGs, deliver underperforming tensorized architectures, suggesting that bigger models, while carrying more expressively, are harder to train, a behavior also noted by Liu et al. (2023a). This is particularly evident looking at the trend of PD-based architectures on FashionMnist.
@@
In Table 2, we compare our best performing architectures, with other state-of-the-art probabilistic models even outside the circuit literature. Our architectures deliver close-to state-of-the-art results, outperforming some VAE- and flow-based models. When compared with RGs learned from data, as it is the case for HCLT, we note that our simpler, data-agnostic alternatives, QTs and QGs, perform equally well or better. Using them instead of the a CL RG saves the quadratic cost needed to learn the corresponding Chow-Liu tree (Dang et al., 2021).
@@
As for the choice of type of sum-product layer, Tucker and CP layers deliver very similar performance on PD. We conjecture that this is due to PD being harder to train in general, as for other RGs the trend changes. In fact, with QT and QG, we observe that Tucker delivers the best bpds for the smallest values for $K$. Scaling it to larger $K$s is impractical however. CP and variants not only scale better (see RQ1 and Fig. 16), but are able to deliver the best bpds for larger $K$. As expected, CP consistently outperforms CPS having more learnable parameters. However, if one has to privilege time over accuracy, CPS can be a useful alternative. Finally, we report results for CPXS layers and learnable mixing layers in Appendix E, along with the results showed in Fig. 17 in tabular form. We confirm that CPS and CPXS layers are equivalently accurate and that one does not have to learn mixing layer parameters in tensorized PCs with DAG-shaped RGs Fig. 11. All these conclusions carry over also to a larger image dataset such as CelebA, using or not the lossless YCoCg color-coding.
@@
Takeaway 2. Our recommendation for a go-to architecture is QG-CP-$K$, with the largest possible $K$ one can squeeze in their GPU memory. If computational resources are not enough, one can trade-off accuracy with speed and use QT-{CP,CPS }-$K$. As a general trend, the simpler the architecture the easier training and scaling are.
@@
For our last research question, consider the problem when a trained circuit with Tucker layers is given, and we want to compress it into a smaller one using CP layers by using our compression pipeline as illustrated in Fig. 12(a) and Fig. 12(b). With this in mind, we investigate the change in performance, if any, w.r.t. the number of tunable parameters. Specifically, for each folded Tucker layer (Eq. Tucker-folded) in the given circuit, parameterized with a tensor $\bm{\mathcal{W}}$ of shape $F{\times}K{\times}K{\times}K$ we compress each tensor slice $\bm{\mathcal{W}}_{f:::}$ by performing non-negative (NN) CP factorization via alternating least squares (Shashua & Hazan, 2005). This optimization eventually delivers a tensor $\bm{\mathcal{W}}^{\prime}$ of shape $F{\times}3{\times}R{\times}K$ for a $R$-ranked factorization.
@@
Figure 18: Compressing Tucker layers into CP layers (Fig. 12(a) $\rightarrow$ Fig. 12(b)) can yield smaller and accurate models as seen when we performing non-negative (NN) CP factorization via alternating least squares (Shashua & Hazan, 2005). In each plot, we report the bpd of a pre-trained Tucker-layered PC (dashed blue line), whose RG, size $K$ and dataset on which it was trained on are detailed on the top. We report the bpd of several $R$-ranked NN-CP factorizations of such PCs (red curves), which we then use as initialization for further fine-tuning (green curves). Finally, we report the bpd of Tucker-compressed PCs (Fig. 12(b)) trained from a random initialization of their parameters (yellow curves).
@@
We sketch the results of our investigation in Fig. 18. As expected, taking a pre-trained Tucker-layered PC (blue dashed line) and compressing its parameters via NN-CP factorization leads to a similar-performing model as the rank $R$ of the approximation increases, as shown by the bpd trend of the red curves in Fig. 18. Interestingly, we observe a key difference between the two region graphs utilized. For tensorized PCs based on PD region graphs, even a rank 1 approximation (i.e. $R=1$) leads to a relatively small bpd lose, while this is not the case for PCs built out of QGs. We conjecture that PD-based PCs have parameter tensors that are of much lower rank than QG-based PCs, and that very deep PCs learn low-rank parameter matrices.
@@
Next, we investigate whether we can use these compressed models as an effective initialization scheme for smaller circuits, which we further train (fine-tune) to maximize the training data likelihood (Eq. 20). Again, we see a different trend when comparing w.r.t. the region graph used, as shown by the bpds encoded as green curves in Fig. 18. Specifically, for PD-based PCs such fine-tuning leads to a quick overfitting already in the first optimization steps, leading to much higher bpds on test data. In contrast, fine-tuning QG-based PCs leads to models that consistently match or even outperform the original Tucker-based PCs (blue dashed line), i.e., we observe green curves consistently being below red curves and breaking the dashed blue lines.
@@
As an additional baseline, we use the architecture of these compressed models (Fig. 12(b)) but train them from scratch: starting from a random initialization of its parameters. Fig. 18 illustrates that the NN-CP initialization can be better than a random one as it leads to better performing models when using QG RGs (yellow curves over green curves). This trend flips when using PD region graphs (yellow curves below green curves), again signaling that much information for these models could be stored in the RG more than in the parameters of the circuit. This, in turn, suggests that while new hierarchical factorizations with highly intricated RGs but very low-rank inner tensors are possible, they might be harder to learn effectively.

### v2

@@ 5.2 Parameter Sharing by Tensor Factorizations
6 Empirical Evaluation: Which RG and Layers to use?
Destructuring modern PC architectures (as well as tensor factorizations) into our pipeline (Fig. 9) allows us to create new tensorized architectures by simply following a mix & match approach (Table 1). At the same time, it helps us understand what really matters between different model classes from the point of views of expressiveness, speed of inference and ease of optimization. We can now in fact easily disentangle key ingredients such as the role of RGs and the choice of composite layers in modern circuit architectures, and pinpoint which is responsible for a boost in performance. For example, HCLTs have been considered as one of the best performing circuit model architectures in recent benchmarks (Liu et al., 2022; 2023a), but until now it has not been clear why they were outperforming other architectures such as RAT-SPNs and EiNets. Within our framework, we can try to answer that question by answering more precise questions: is it the effect of their RG that is learned from data (Section 4.1)?, the use of their composite sum-product layer parameterization (Section 5.1)? or are other hyperparameter choices responsible? (spoiler: it is going to be the use of CP layers).
Specifically, in this section we are interested in answering the following three research questions following a rigorous empirical investigation. RQ1) What are the computational resources needed (time and GPU memory) at test and training time for some of the many tensorized architectures we can now build? RQ2) What is the impact of the choice of RG and composite sum-product layer on the performance of tensorized circuits trained as distribution estimators? RQ3) Can we retain (most of) the performances of pre-trained tensorized PCs using Tucker layers if we factorize these into CP layers as illustrated in Fig. 13(a) $\rightarrow$ Fig. 13(b)? Note that we are not asking what is the impact of folding (Section 4.4), as we already know the answer: folding is essential for large-scale tensorized architectures. As such, throughout all experiments, we use folded tensorized circuits. We emphasise that the aim of our experiments is not to reach state-of-the-art results in distribution estimation, but rather to understand the role of the ingredients of tensorized circuit architectures. All experiments were run on a single NVIDIA RTX A6000 GPU with 48GB of memory. Our code is available at github.com/april-tools/uni-circ-le.
@@ 6 Empirical Evaluation: Which RG and Layers to use?
We will mainly evaluate our architectures by performing distribution estimation on image datasets. We use the Mnist-family, which includes 6 datasets of gray-scale $28\times 28$ images—Mnist (LeCun et al., 2010), FashionMnist (Xiao et al., 2017), and EMNIST with its 4 splits (Cohen et al., 2017)—and the CelebA dataset down-scaled to $64\times 64$ (Liu et al., 2015), which we explore in two versions: one with RGB pixels and the other with pixels preprocessed by the lossless YCoCg color-coding (Malvar & Sullivan, 2003), as recent results suggested that such a transform can greatly lower bpds. [^8] We take this evidence from Liu et al. (2023a; b), which use however a lossy variant of the YCoCg transform that unfortunately artificially inflates likelihoods. As such, their bpds for PCs are not directly comparable with ours, nor with the other deep generative models in their tables. We confirmed this issue in their evaluation protocol via personal communication. Furthermore, we perform experiments on tabular data with continuous variables. In particular, we will evaluate different tensorized layers by performing density estimation on 5 UCI datasets, as they are typically used to evaluate normalizing flows (Papamakarios et al., 2017). We report the statistics of the UCI dataset in Table E.5.
@@
We train circuits to estimate the probability distribution that is assumed to have generated the images, considering each pixel as a random variable. As such, the input units in the circuit represent Categorical distributions having 256 values. For RGB images, we associate three Categorical distribution units per pixel (one per color channel). Instead, for the 5 UCI datasets (Table E.5), we use input units representing univariate Gaussian distributions, and we learn both the means and the standard deviations. We perform maximum likelihood by stochastic gradient ascent, i.e., want to maximize the following objective
@@
where $Z=\sum_{\bm{\mathrm{x}}}c(\bm{\mathrm{x}})$ is the partition function of the PC $c$ [^9] After training, one can efficiently “embed” the normalization constant in the parameters of a PC, effectively renormalizing them (and thus yielding a partition function $Z$ equal to $1$), as detailed by Peharz et al. (2015)., and $\mathcal{B}$ a batch of training data. After some preliminary experiments, we found that optimizing PCs with Adam (Kingma & Ba, 2015) using a learning rate of $10^{-2}$ delivered, on average, the best performing models for the datasets we considered. We also settled to reparameterize the circuit sum parameters via clamping and setting $\epsilon=10^{-19}$ (Eq. 9) after each optimization step as to keep them non-negative, as it was giving the best learning dynamics among all possible reparameterizations (Section 3.2). In the following, we will summarize our findings when answering RQ1-3, while distilling recommendations for practitioners on how to build and learn circuits.
@@
For these experiments, we consider the following RGs: PD, as commonly used in architectures such as RAT-SPNs and EiNets, and the two novel light-weight and data-agnostic RGs we introduced in Section 4.1, QTs  [^10] Throughout our experiments, we will refer to QT-4 simply as QT. and QGs . We do not consider RND as it is usually just a balanced binary tree (Peharz et al., 2020c), and as such would yield the same time and memory performance of a QT. For the same reason we do not consider CL as they are tree RGs that end up being quasi-balanced after being rooted. [^11] The root is chosen to be the barycenter of the graph to increase parallelism (Dang et al., 2021; 2022c). For layers, we consider Tucker (Eq. Tucker-layer), CP (Eq. CP-layer), CPS (Eq. CPS-layer) and CPXS (Section 5.2).
@@
We emphasise that non-folded versions of these architectures, e.g. RAT-SPNs (Peharz et al., 2020c), can be orders of magnitude slower, hindering both learning and deployment in practice. In Fig. E.1, we show the results of the same benchmark reported in Fig. 17 but for the CelebA dataset, which is more challenging because it is equivalent to perform distribution estimation on a much higher dimensional space ($12,288=64\times 64\times 3$ instead of $784=28\times 28\times 1$). [^12] Note that for our RQ1, all image datasets with the same resolutions would yield the very same results. From this additional experiment, we conclude that even in higher dimensions the scaling trend of RGs and layers is the same. Finally, in Fig. E.3, we zoom on a comparison between CPS and CPXS. There, we show that for the same choice of RG and $K$, CPS and CPXS layers require the same time/space resources as expected, with CPXS only being slightly faster at training-time.
Takeaway 1. QT and QG should be preferred to PD as RGs if we want to scale circuits, with the former being more scalable than the latter. Layer-wise, CP layers scale, as expected, to larger values of $K$ than Tucker layers and for even larger layers parameter sharing (CPS, CPXS) is recommended.
@@
We now test our tensorized PCs as distribution estimators and we consider our mixed&matched architecture from RQ1. For each architecture, we vary the model size by varying $K$, the number of units for each layer, in $\{16,32,64,128,256,512\}$ for the MNIST family and up to $256$ for CelebA. To assess the effect of learning RGs from data, we compare against HCLTs (CL-$\text{CP}^{\top}$in our nomenclature) as reported by Dang et al. (2022a). We use a batch size of 256, and train for at most 200 epochs stopping training if the validation log-likelihood does not improve after 5 epochs. We use the average test-set bits-per-dimension (bpd) as the evalutation criterion, i.e. $\mathsf{bpd}(\mathcal{D},c)=-\mathcal{L}(\mathcal{D},c)/(d\cdot\log 2)$, where $d$ is the number of features in dataset $\mathcal{D}$ and $\mathcal{L}$ is defined as in Eq. 18.
@@
In Fig. 18 we report the average test-set bpd on Mnist, FashionMnist and CelebA. An immediate visible pattern emerges when comparing the architectures w.r.t. the choice of RG: Both QT- and QG-based architectures outperform those based on PD, and also manage to scale to larger datasets like CelebA. On average, the best performing architectures are those built out of QGs. This is expected as such RGs, different from QTs, allow different partitionings for a same region (and therefore require the usage of mixing layers as discussed in Eq. Mixing-layer). The PD region graphs, despite being DAG-shaped as QGs, deliver underperforming tensorized architectures, suggesting that bigger models, while being more expressive, are harder to train, a behavior also noted by Liu et al. (2023a). This is particularly evident looking at the trend of PD-based architectures on FashionMnist.
@@
In Table 2, we compare our best performing architectures, with other state-of-the-art models even outside the circuit literature. Our architectures deliver close-to state-of-the-art results, and outperforms some VAE- and flow-based models. When compared with RGs learned from data, as it is the case for HCLT, we note that our simpler, data-agnostic alternatives, QTs and QGs, perform equally well or better. Using them instead of HCLTs avoids the quadratic cost to learn the corresponding Chow-Liu tree (Dang et al., 2021).
@@
As for the choice of type of sum-product layer, Tucker and CP layers deliver very similar performance on PD. We conjecture that this is due to PD being harder to train in general, as for other RGs the trend changes. In fact, with QT and QG, we observe that Tucker delivers the best bpds for the smallest values for $K$. Scaling it to larger $K$s is impractical however. CP and variants not only scale better (see RQ1 and Fig. 17), but are able to deliver the best bpds for larger $K$. As expected, CP consistently outperforms CPS having more learnable parameters. However, if one has to privilege time over accuracy, CPS can be a useful alternative. Finally, we report results for CPXS layers and learnable mixing layers in Appendix E, along with the results showed in Fig. 18 in tabular form. We confirm that CPS and CPXS layers are equivalently accurate and that one does not have to learn mixing layer parameters in tensorized PCs with DAG-shaped RGs (Fig. 12). All these conclusions carry over also to a larger image dataset such as CelebA, with or without the lossless YCoCg color-coding.
Density estimation on tabular datasets.
Finally, we perform density estimation experiments on UCI datasets (Table E.5), and compare the results achieved by tensorized PCs constructed by our pipeline by parameterizing a RND RG (Section 4.1) with either CP or Tucker layers. To give context to our results, we show the average test log-likelihoods achieved by normalizing flow models (Papamakarios et al., 2021) that are often evaluated on UCI datasets: MADE (Germain et al., 2015), RealNVP (Dinh et al., 2017), MAF (Papamakarios et al., 2017) and NSF (Durkan et al., 2019). As additional baselines, we show results of other PCs supporting tractable marginalization: a single multivariate Gaussian, Einsum networks (Peharz et al., 2020a) with input layers encoding flows (EiNet-LRS) (Sidheekh et al., 2023), and TTDE (Novikov et al., 2021). We emphasize that both EiNet-LRS and TTDE can be built using our pipeline and characterized with our nomenclature, the former as RND RGs parameterized by Tucker layers, and the latter as LT RGs parameterized by $\text{CP}^{\top}$layers (Table 1). Table 3 shows that CP layers to deliver better performances than Tucker layers on high-dimensional UCI datasets and therefore in the case of deeper tensorized PCs. On the other hand, Tucker layers outperform CP layers on the lower-dimensional UCI datasets. We believe this is due the parameters of Tucker layers being more difficult to train and scale, similarly to our observation for Mnist and FashionMnist in the case of QG RGs in Fig. 18. We further detail in Section E.1 the experimental setting, and show in Fig. E.5 the results achieved by varying the layer width $K$.
Table 3: Tucker layers are harder to scale than CP layers on high-dimensional UCI datasets. We show the best average test log-likelihoods achieved by normalizing flow models (top) and tensorized PCs that can be instantiated from our pipeline (bottom). See main text for their description. Tensorized PCs obtained by parameterizing random binary tree RGs (Section 4.1) with CP (RND-CP) perform better on higher-dimensional datasets Hepmass and MiniBooNE than those with Tucker layers (RND-Tucker), while the latter have an advantage on lower-dimensional datasets such as Power and Gas. For RND-CP and RND-Tucker, we report the layer width ($K$) of the best performing model as a subscript of the log-likelihoods. Fig. E.5 shows training and test log-likelihoods achieved by varying the layer width $K$. Details in Section E.1.
Power Gas Hepmass M.BooNE BSDS300
MADE (Germain et al., 2015) -3.08 3.56 -20.98 -15.59 148.85
RealNVP (Dinh et al., 2017) 0.17 8.33 -18.71 -13.84 153.28
MAF (Papamakarios et al., 2017) 0.24 10.08 -17.73 -12.24 154.93
NSF (Durkan et al., 2019) 0.66 13.09 -14.01 -9.22 157.31
Gaussian -7.74 -3.58 -27.93 -37.24 96.67
EiNet-LRS (Sidheekh et al., 2023, RND-Tucker) 0.36 4.79 -22.46 -34.21
TTDE Novikov et al. (2021, LT-$\text{CP}^{\top}$) 0.46 8.93 -21.34 -28.77 143.30
RND-CP $0.28_{256}$ $5.01_{256}$ $-22.52_{64}$ $-30.69_{128}$ $120.82_{64}$
RND-Tucker $0.52_{64}$ $8.41_{256}$ $-23.47_{32}$ $-31.30_{8}$ $119.09_{64}$
@@
Takeaway 2. In the case of image datasets, our recommendation for a go-to architecture is QG-CP-$K$, with the largest possible $K$ one can squeeze in their GPU memory. If computational resources are not enough, one can trade-off accuracy with speed and use QT-{CP,CPS}-$K$. As a general trend, the simpler the architecture the easier training and scaling are. This is also suggested by our results on UCI datasets, where the simpler CP layers can perform better for high-dimensional datasets than Tucker.
@@
For our last research question, we consider the problem when a trained circuit with Tucker layers is given, and we want to compress it into a smaller one using CP layers by using our compression pipeline as illustrated in Fig. 13(a) and Fig. 13(b). With this in mind, we investigate the change in performance, if any, w.r.t. the number of tunable parameters. Specifically, for each folded Tucker layer (Eq. Tucker-folded) in the given circuit, parameterized with a tensor $\bm{\mathcal{W}}$ of shape $F{\times}K{\times}K{\times}K$ we compress each tensor slice $\bm{\mathcal{W}}_{f:::}$ by performing non-negative (NN) CP factorization via alternating least squares (Shashua & Hazan, 2005). This optimization eventually delivers a tensor $\bm{\mathcal{W}}^{\prime}$ of shape $F{\times}3{\times}R{\times}K$ for a $R$-ranked factorization.
@@
Figure 19: Compressing Tucker layers into CP layers (Fig. 13(a) $\rightarrow$ Fig. 13(b)) can yield smaller and accurate models as seen when we performing non-negative (NN) CP factorization via alternating least squares (Shashua & Hazan, 2005). In each plot, we report the bpd of a pre-trained Tucker-layered PC (dashed blue line), whose RG, size $K$ and dataset on which it was trained on are detailed at the top. We report the bpd of several $R$-ranked NN-CP factorizations of such PCs (red curves), which we then use as initialization for further fine-tuning (green curves). Finally, we report the bpd of Tucker-compressed PCs (Fig. 13(b)) trained from a random initialization of their parameters (yellow curves).
@@
We sketch the results of our investigation in Fig. 19. As expected, taking a pre-trained Tucker-layered PC (blue dashed line) and compressing its parameters via NN-CP factorization leads to a similar-performing model as the rank $R$ of the approximation increases, as shown by the bpd trend of the red curves in Fig. 19. Interestingly, we observe a key difference between the two region graphs utilized. For tensorized PCs based on PD region graphs, even a rank 1 approximation (i.e. $R=1$) leads to a relatively small bpd loss, while this is not the case for PCs built out of QGs. We conjecture that PD-based PCs have parameter tensors that are of much lower rank than QG-based PCs, and that very deep PCs learn low-rank parameter matrices.
@@
Next, we investigate whether we can use these compressed models as an effective initialization scheme for smaller circuits, which we further train (fine-tune) to maximize the training data likelihood (Eq. 18). Again, we see a different trend when comparing w.r.t. the region graph used, as shown by the bpds encoded as green curves in Fig. 19. Specifically, for PD-based PCs such fine-tuning leads to a quick overfitting already in the first optimization steps, leading to much higher bpds on test data. In contrast, fine-tuning QG-based PCs leads to models that consistently match or even outperform the original Tucker-based PCs (blue dashed line), i.e., we observe green curves consistently being below red curves and crossing the dashed blue lines.
@@
As an additional baseline, we use the architecture of these compressed models (Fig. 13(b)) but train them from scratch: starting from a random initialization of its parameters. Fig. 19 illustrates that the NN-CP initialization can be better than a random one as it leads to better performing models when using QG RGs (yellow curves over green curves). This trend flips when using PD region graphs (yellow curves below green curves), again signaling that much information for these models could be stored in the RG rather than in the parameters of the circuit. This, in turn, suggests that while new hierarchical factorizations with highly intricate RGs but very low-rank inner tensors are possible, they might be harder to learn effectively.
@@
7 Additional Related Work
Tensor networks and PGMs. TNs (Orús, 2013) are widely used to model many-body systems in physics and quantum mechanics (Schollwoeck, 2010), and have been used to simulate quantum computations on classical hardware (Markov & Shi, 2008). They have been applied more recently for machine learning applications (Stoudenmire & Schwab, 2016; Han et al., 2018; Efthymiou et al., 2019; Bonnevie & Schmidt, 2021). As they essentially an alternative formalism for probabilistic graphical models over discrete variables (Koller & Friedman, 2009), people have started drawing connections between the two formalisms. For example, Bonnevie & Schmidt (2021) connects non-negative MPS/TTs to PGMs and offers routines for probabilistic reasoning. Similarly, Glasser et al. (2020) explores the same connection, but instead of drawing TNs as PGMs, they draw them as factor graphs (Kschischang et al., 2001).
Instead, by representing non-negative tensor factorizations as monotonic PCs, we effortlessly unlock the developed theory and algorithms required to perform complex probabilistic inference, with possible applications in lossless compression (Liu et al., 2022), neuro-symbolic AI with correctness guarantees (Ahmed et al., 2022) and constrained text generation (Zhang et al., 2023). Moreover, results about the succinctness or expressive efficiency of these factorizations (Glasser et al., 2019) have been used recently to prove circuit complexity lowerbounds (Loconte et al., 2024; 2025). Finally, Loconte & Vergari (2025) took inspiration from canonical forms in tensor networks (Schollwoeck, 2010; Bonnevie & Schmidt, 2021) to parameterize already-normalized squared circuits generalizing squared MPS/TTs and TTNs (Section 2.4), and to devise a more efficient marginalization algorithm within the circuit language.
Probabilistic circuits and PGMs. The modern formulation of PCs has been introduced for the first time in (Vergari et al., 2019b) as a unifying framework for several existing tractable probabilistic models (TPMs) including arithmetic circuits, (Darwiche, 2001), probabilistic decision graphs (Jaeger, 2004), and-or graphs (Marinescu & Dechter, 2009), cutset networks (Rahman et al., 2014), sum-product networks (Poon & Domingos, 2011) and more (Choi et al., 2020). The aim of PCs has been to abstract away from the different syntaxes and model formalisms of the above TPMs and focus on structural properties that enable tractable inference in each. Non-negative tensor factorizations and tensor networks have been underlooked in this effort so far. Several ways to compile discrete PGMs into PCs (or one of the above formalisms) have been devised in the past (Oztok & Darwiche, 2017; Shen et al., 2016; Choi et al., 2013). These compilation techniques yield sparse deterministic circuits, and only recently PCs have started to be represented first in code (Peharz et al., 2020c; a; Liu & Van den Broeck, 2021b) and then formally (Loconte et al., 2024) as tensorized architectures. Perhaps this lack of tensorized compilation targets has hidden the connection between PCs and matrix and tensor factorizations. The closest connection we are aware of can be found in Jaini et al. (2018b): they bridge sum-product networks to hierarchical mixture models and HMMs and hint at a connection with tensorial mixture models (Sharir et al., 2017) a variant of hierarchical Tucker (Def. 5).
Matrix factorizations, circuit complexity, and tensor networks expressiveness. Finding lower bounds to the rank of matrix factorizations can be used as a proxy to prove lower bounds to the size of circuits satisfying particular structural properties (de Colnet & Mengel, 2021). Proving an exponential (w.r.t. the number of variables) size lower bound for a class of circuits shows a limitation on which functions they can compute in polynomial time and number of parameters, thus allowing us to precisely separate circuit classes in terms of their expressiveness (Valiant, 1979; Martens & Medabalimi, 2014). Recently, lower bounding the non-negative rank (Gillis, 2020) and the square root rank (Fawzi et al., 2014; Lee & Wei, 2014) of matrices has been used to draw an expressiveness hierarchy of classes of PCs with negative real and complex-valued parameters for distribution estimation (Loconte et al., 2024; 2025). Since circuits generalize many tensor network factorizations (see Section 2.4), showing size lower bounds for a class of circuits can be used to show size lower bounds for tensor networks regardless of their structure, e.g. as shown by Loconte et al. (2025) in generalizing a known rank lower bound for Born machines obtained by squaring a MPS/TT with real-valued tensors (Glasser et al., 2019).